### v1

@@ 1 Introduction
It has been recently shown that one can obtain finite approximations via state and action quantization for Markov decision processes (MDPs) with uncountable Polish spaces, as well as run both empirical model learning and Q-learning to arrive at near optimal solutions (see.e.g. [24, 27]). Such studies have primarily focused on the case with compact spaces and uniform quantization and with only asymptotic convergence results for the case with non-compact spaces. To this end, the goal of this paper is to design quantizers in such a finite model approximation and learning framework for continuous space MDPs. Notably, this paper develops a mathematical framework and formulation paving the way for non-uniform quantizer design and associated performance bounds, which are then also applicable to MDPs with non-compact spaces.
@@ 1.1 Related Literature
For stochastic control problems with continuous state and action spaces, approximations are inevitable for computational methods. A common approach in reinforcement learning is function approximation, where the value function (Q-function) of the control problem is approximated using a parametrized family of functions. Convergence of policy evaluation methods are known under linear function approximation where the parametrized family of functions are formed by the span of finitely many basis functions [33]. However, learning optimal Q-functions with linear function approximation is known to be unstable in general [2] except in special cases. For general linear function approximation, [20] has shown that under certain a class of exploration policies, optimal Q-value learning with linear function approximation remains bounded. The special cases where the convergence can be guaranteed include when (i) the exploration policy is already close to the greedy policy of the learning iterations [18], and (ii) the stage-wise cost function, the transition model, and thus the optimal Q-functions are perfectly represented by the basis functions, i.e. they belong to the span of the basis functions [10, 23]. These assumptions, however, are restrictive in general, since it is unrealistic to assume near-optimal exploration or perfect linear representability.
A particularly powerful, though computationally demanding, special case is when the basis functions are indicator functions of quantization bins, which form an orthonormal basis, and the convergence analysis can mathematically be justified. In particular, [16] showed that the Q-values learned under quantization based learning correspond to an auxiliary finite control problem, which is a finite approximation of the original control problem with a particular weight measure on the quantization bins. This observation has an important consequence that the model-based approaches using state space quantization and the quantized Q-learning coincide and can be used interchangeably, even under very mild continuity conditions on the transition kernels [27].
@@
Another direction of research with provable loss guarantees is kernel-based methods [22, 21, 3, 6, 35, 7] where the cost and the transition models are estimated via empirical kernel regression or otherwise simulations are used to update the Bellman optimality equations. One can then find the control policies based on the learned models using model-based approaches.
Most prior work focuses on approximation of control problems with compact and bounded spaces. There is a limited number of provable guarantees for the approximation of control problems with unbounded state spaces. In [27] asymptotic near optimality of the quantization based approximations was established for non-compact state spaces as the quantization rate grows to infinity. However, there is no convergence rate guarantees for this approach in terms of expected quantization error on the state space and learning was not studied in this context.
@@ 1.2 Model and Cost Criteria
where $\mathcal{T}(\cdot|x,u)$ is a stochastic kernel (that is, a regular conditional probability measure) from $\mathds{X}\times\mathds{U}$ to $\mathds{X}$, $\mathcal{B}(\mathds{X})$ is the Borel $\sigma$-algebra of $\mathds{X}$, and $(X,U)_{[0,t-1]}$ is the set of state-action pairs up until t1t-1.
We will be interested in the following performance criteria: The first one is the infinite-horizon discounted expected cost
@@ 1.3 Contributions
In this paper, we make the following contributions:
  • (i)
    In Theorem 2, we refine and computationally improve the upper bounds given in [16] which involve admissible policies to ones that only involve stationary policies. This facilitates an analysis involving occupation measures as well as invariant measures for discounted and ergodic cost criteria, respectively; which is then utilized later in the paper.
  • (ii)
    In Theorems 3 9, we derive bounds in terms of occupation measures, to represent the loss in terms of occupation and invariant measures, respectively for the discounted and average cost criteria. In Corollary 6, we optimize the quantization design by choosing the representative points as medians (in the $\ell_{1}$ sense) with respect to the occupation measures and arrive at an explicit error bound. We thus provide a convergence rate analysis of quantization based approximations which is convenient for stochastic analysis and which is also applicable for non-compact state spaces.
  • (iii)
    In Theorems 7 and 8 under Foster-Lyapunov conditions, we derive explicit error bounds which decay to zero as the quantization gets finer and show better performance than uniform quantization in MDPs with non-compact state spaces under discounted cost criteria. We extend our analysis to the average cost criterion by obtaining explicit error bounds in Theorem 11.
  • (iv)
    In Theorem 14, we extend our analysis on quantizer design to quantized Q-learning and empirical model learning (which are equivalent in performance). We obtain the error bounds which depend, unlike the planning problem above, on the invariant measure induced by the exploration policy in Theorem 13. For non-compact spaces, in Theorem 14, we show that the approximation error diminishes as quantization becomes finer under Foster–Lyapunov conditions, despite constraints on the dependence of the weighting measures on the exploration policy and quantization bins, and we derive explicit error bounds.
@@ 2.2 Refined Error Bounds for the Discounted Cost Criterion
The above ensures that optimal policies exist. Furthermore, by [24, Lemma 3.19] and [24, Theorem 3.16] (see also [26]), any MDP with a weakly continuous transition probability kernel can be approximated by an MDP with finite action spaces. Accordingly, in the sequel, we assume that $\mathds{U}$ is finite.
Remark 2.2.
[…]
@@
where the probability measure $\mathbb{P}_{x_{0}}^{\gamma_{s}}$ over the state and action process is defined by the initial condition $x_{0}$, the stationary policy $\gamma_{s}$, and the transition kernel $\mathcal{T}$.
For a product set $A\times\mathds{U}$ with $A\subseteq\mathds{X}$ measurable we obtain
@@
To minimize the error bound obtained, we optimize the quantizer by choosing the centroids of the quantization bins properly. The following lemma will be useful in choosing the centroids.
@@
[5] Let $\mathds{X}\subseteq\mathds{R}$, and let $X$ be a random variable with distribution $\mu$ over a quantization bin $B_{i}$. Suppose that $\mathbb{E}(|X|)<\infty$. Then, the value $y_{i}$ that minimizes the expected distortion $\mathbb{E}[|X-y_{i}|]$ is the median of $X$ with respect to the probability measure $\mu$ over $B_{i}$ .
For a proof, see [5].
According to Lemma 4, we establish that for a given quantization bin $B_{i}$, the expected distortion is minimized when the representative point $y_{i}$ is chosen as the median of $B_{i}$ with respect to the probability measure $\mu$ over $B_{i}$. This turns the normalized measure $\hat{\pi}_{y_{i}}$ into the Dirac measure $\delta_{y_{i}}$ concentrated at $y_{i}$. This result can be also applied to higher dimensional state spaces in $\mathbb{R}^{n}$ according to the following theorem.
The result can be extended to higher dimensions when considering the $\ell_{1}$ norm in $\mathbb{R}^{n}$.
Lemma 5.
Let $X=(X_{1},X_{2},\dots,X_{n})$ be a random vector with finite expected absolute deviations in each coordinate. Then, the point $y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*})$ that minimizes $\mathbb{E}[\|X-y_{i}\|_{1}]$ is obtained by choosing each $y_{i,k}^{*}$ to be a median of the marginal distribution of $X_{k}$ over the bin $B_{i}$, for $k=1,2,\dots,n$.
The proof of the above is immediate: Let $X$ be an n -dimensional random vector, and $y_{i}\in\mathbb{R}^{n}$ . The expected $\ell_{1}$ distortion is:
@@
where $\mu_{k}$ is the marginal distribution of $X_{k}$ over the bin $B_{i}$. Since the sum of convex functions is convex, $\mathbb{E}[\|X-y_{i}\|_{1}]$ is convex in $y_{i}$ . Each component $y_{i,k}$ appears only in the term involving $x_{k}$, so we can minimize each term separately. By the same reasoning as in the scalar case, as in Lemma 4, $\int_{B_{i}}|x_{k}-y_{i,k}|\mu_{k}(dx)$ is minimized when $y_{i,k}$ is a median of $X_{k}$ with respect to the probablity measure $\mu_{k}$ over $B_{i}$. Therefore, the point $y^{*}_{i}$ that minimizes the expected $\ell_{1}$ distortion is found as:
$y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*}),\quad\text{for}\quad k=1,2,\dots,n$ (8)
where each $y_{i,k}^{*}$ is the median of the quantization bin $B_{i}$ with respect to the marginal distribution $\mu_{k}$.
@@
Assume that the state space $\mathds{X}\subseteq\mathbb{R}$ and let $b\geq 0$, $\alpha>0$ and $V:\mathds{X}\to[0,\infty)$. Assume the state process $\{X_{t}\}$ satisfies the following condition:
$\mathbb{E}^{\gamma_{s}}[V(X_{t+1})|X_{t}=x,U_{t}=u]\leq V(x)-\alpha V(x)+b,\quad x\in\mathds{X},u\in\mathds{U}$ (9)
for all $\gamma_{s}\in\Gamma_{s}$ where $V(x)=|x|^{m}$, for some $m>1$.
Then, under Assumption 2.1 and Assumption 2.2, with $M$ denoting the number of bins, there exists a quantizer which leads to the following error bound for any initial state $x_{0}\in\mathds{X}$:
$C:=\frac{|x_{0}|^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}$
@@
The next result extends Theorem 7 to an nn–dimensional state space $\mathds{X}\subseteq\mathbb{R}^{n}$ (with the $\ell_{1}$–norm).
Theorem 8.
Let $b\geq 0$, $\alpha>0$, and define the Lyapunov function $V(x)=\|x\|_{1}^{m}$ with exponent $m>n$. Assume the controlled process $\{X_{t}\}$ satisfies the drift condition
$\mathbb{E}^{\gamma_{s}}\bigl[V(X_{t+1})\mid X_{t}=x,U_{t}=u\bigr]\leq V(x)-\alpha V(x)+b,\qquad x\in\mathds{X},u\in\mathds{U}$
for all $\gamma^{s}\in\Gamma_{s}$. Under Assumption 2.1 and Assumption 2.2, let $M$ be the total number of hyper-cubic bins in the uniform quantizer. Then, for every initial state $x_{0}\in\mathds{X}$ we have that,
$\bigl|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\bigr|\leq\Bigl(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\Bigr)\frac{(2n+1)C^{1/m}}{M^{\frac{1}{n}-\frac{1}{m}}(1-\beta)},$ (10)
where
$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
Proof.
The proof is in Appendix D. ∎
The explicit quantizer is presented in the proof of Theorem 8.
@@ 2.3 Refined Error Bounds for the Average Cost Criterion
The proof is in Appendix E
Now, we state the following theorem, which will be useful in the proof of our next result:
Theorem 10 (Comparison Theorem).
[19, Theorem 14.2.2] Let $V:\mathds{X}\to[0,\infty)$ and $f,g:\mathds{X}\to[0,\infty)$. Let $\{x_{n}\}$ be a Markov chain on $\mathds{X}$. If the following condition is satisfied:
$\int_{\mathds{X}}P(x,dy)V(y)\leq V(x)-f(x)+g(x),\quad x\in\mathds{X},$
then, for any stopping time $\tau$ with $\mathbb{P}(\tau<\infty)=1$, it holds that
$\mathbb{E}\left[\sum_{t=0}^{\tau-1}f(x_{t})\right]\leq V(x_{0})+\mathbb{E}\left[\sum_{t=0}^{\tau-1}g(x_{t})\right].$
Proof.
See [34]. ∎
Now, we present our result under Lyapunov condition similar to the quantizer design for discounted cost criterion.
@@
Assume $\mathds{X}\subseteq\mathbb{R}$, $f(x)=\|x\|^{m}$ with $m>1$, and suppose
$\sup_{x,u}\mathbb{E}[V(X_{t+1})\mid X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$
@@
$|\hat{J}_{avg}(x_{0})-J_{avg}^{*}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{3b^{1/m}}{M^{1-1/m}}.$ (13)
@@ 3.1 Quantized Q-Learning Algorithm and its Convergence
As noted in [14], quantizing the state space of a continuous MDP converts the problem into a Partially Observable Markov Decision Process (POMDP) which is then non-Markovian. In particular, with the quantization process, the controller has partial observation. Define
$Y_{t}=q(X_{t}),$
where $q:\mathbb{X}\rightarrow\mathbb{Y}$ is the quantizer, which maps the continuous state space $\mathbb{X}$ to the finite set of representative points $\mathbb{Y}=\{y_{1},y_{2},\ldots,y_{M}\}$.
The observation model in our setup is described as:
$O(y_{i}\mid x)=\mathbb{P}(Y_{t}=y_{i}\mid X_{t}=x)=\mathds{1}_{B_{i}}(x),$
where $\mathds{1}_{B_{i}}(x)$ is the indicator function. In this model, the agent has access to only the representative points $Y_{t}=y_{i}$, which indicates that the true state $X_{t}$ resides in the quantization bin $B_{i}$. In other words, the true state $X_{t}$ remains hidden from the agent. In the quantized MDP, the agent observes $Y_{t}=q(X_{t})$.
@@
We note that a sufficient condition for the second item above is that the state process $\{X_{t}\}_{t\geq 0}$ is positive Harris recurrent.
The following result is proved by Kara and Yüksel [11]. It states that the algorithm in (1) converges to the optimal Q-function of the approximate belief MDP under Assumption 3.1:
We have the following convergence result:
Theorem 12 (Kara et al., 2023, Theorem 9 [16]).
@@ 3.2 Empirical Model Learning and Equivalence with Quantized Q-Learning
Observe that the above is, see e.g. [15, Theorem 2.1], equal to the almost sure limit of the empirical expression on the right hand side below:
@@
Likewise, by [15, Theorem 2.1], the above is the almost sure empirical limit of of the right hand side below:
@@
An interpretation of the above result then is that one can first obtain the approximate model given with (20-22) by forcing the data into a Markovian model for both the empirical cost estimate (21) and empirical transition kernel estimate (23), and then solve the MDP as if this Markovian empirically constructed model is the actual one, instead of running Q-learning whose limit is then optimal precisely for this learned/empirically constructed model. A benefit of such a model-based approach is that one can have better sample complexity bounds compared with Q-learning for certain applications, see e.g. [35] (see also [13, Section 5.1]).
@@
In the planning framework presented in the previous section, we had the flexibility to select the weighting measures $\pi_{y_{i}}$ over the quantization bins arbitrarily. This allowed us to minimize the expected loss by choosing $\pi_{y_{i}}$ as a Dirac measure centered at the median of each bin. In contrast, within the learning context, the weighting measure $\pi_{y_{i}}$ is dictated by the exploration policy and is inherently dependent on the structure of the quantization bins $\{B_{i}\}$. Specifically, $\pi_{y_{i}}$ corresponds to the normalized invariant distribution over bin $B_{i}$ induced by the exploration policy; that is, it is the conditional probability measure given that the state is in $B_{i}$:
$\pi_{y_{i}}(A)=\frac{1_{\{A\in B_{i}\}}\pi_{\gamma^{*}}(A)}{\pi_{\gamma^{*}}(B_{y})},\qquad A\in B_{i},A\in{\cal B}(\mathbb{X})$

### v2

@@ 1 Introduction
It has been recently shown that one can obtain finite approximations via state and action quantization for Markov decision processes (MDPs) with uncountable Polish spaces, as well as run both empirical model learning and Q-learning to arrive at near optimal solutions (see.e.g. [23, 26]). Such studies have primarily focused on the case with compact spaces and uniform quantization and with only asymptotic convergence results for the case with non-compact spaces. To this end, the goal of this paper is to design quantizers in such a finite model approximation and learning framework for continuous space MDPs.
@@ 1.1 Related Literature
For stochastic control problems with continuous state and action spaces, approximations are inevitable for computational methods. A common approach in reinforcement learning is function approximation, where the value function (Q-function) of the control problem is approximated using a parametrized family of functions. Convergence of policy evaluation methods are known under linear function approximation where the parametrized family of functions are formed by the span of finitely many basis functions [32]. However, learning optimal Q-functions with linear function approximation is known to be unstable in general [2] except in special cases. For general linear function approximation, [19] has shown that under a certain class of exploration policies, optimal Q-value learning with linear function approximation remains bounded. The special cases where the convergence can be guaranteed include when (i) the exploration policy is already close to the greedy policy of the learning iterations [17], and (ii) the stage-wise cost function, the transition model, and thus the optimal Q-functions are perfectly represented by the basis functions, i.e. they belong to the span of the basis functions [10, 22]. These assumptions, however, are restrictive in general, since it is unrealistic to assume near-optimal exploration or perfect linear representability.
A particularly powerful, though computationally demanding, special case is when the basis functions are indicator functions of quantization bins, which form an orthonormal basis, and the convergence analysis can mathematically be justified. In particular, [15] showed that the Q-values learned under quantization based learning correspond to an auxiliary finite control problem, which is a finite approximation of the original control problem with a particular weight measure on the quantization bins. This observation has an important consequence that the model-based approaches using state space quantization and the quantized Q-learning coincide and can be used interchangeably, even under mild continuity conditions on transition kernels [26].
@@
Another direction of research with provable loss guarantees is kernel-based methods [21, 20, 3, 6, 34, 7] where the cost and the transition models are estimated via empirical kernel regression or otherwise simulations are used to update the Bellman optimality equations. One can then find control policies based on the learned models using model-based approaches.
Most prior work focuses on approximation of control problems with compact spaces. There is a limited number of provable guarantees for the approximation of control problems with unbounded state spaces. In [26] asymptotic near optimality of the quantization based approximations was established for non-compact state spaces as the quantization rate grows to infinity. However, there is no convergence rate guarantees for this approach in terms of expected quantization error on the state space and learning was not studied in this context.
@@ 1.2 Model and Cost Criteria
where $\mathcal{T}(\cdot|x,u)$ is a stochastic kernel (that is, a regular conditional probability measure) from $\mathds{X}\times\mathds{U}$ to $\mathds{X}$, $\mathcal{B}(\mathds{X})$ is the Borel $\sigma$-algebra of $\mathds{X}$, and $(X,U)_{[0,t-1]}$ is the set of state-action pairs up until t1t-1. We will be interested in the following performance criteria: The first one is the infinite-horizon discounted expected cost
@@ 1.3 Contributions
In this paper, we make the following contributions: (i) In Theorem 2, we refine and computationally improve the upper bounds given in [15] which involve admissible policies to ones that only involve stationary policies. This facilitates an analysis involving occupation measures as well as invariant measures for discounted and ergodic cost criteria, respectively; which is then utilized later in the paper. (ii) In Theorems 3 7, we derive bounds in terms of occupation measures, to represent the loss in terms of occupation and invariant measures, respectively for the discounted and average cost criteria. In Corollary 5, we optimize the quantization design by choosing the representative points as medians (in the $\ell_{1}$ sense) with respect to the occupation measures and arrive at an explicit error bound. We thus provide a convergence rate analysis of quantization based approximations which is convenient for stochastic analysis and which is also applicable for non-compact state spaces. (iii) In Theorem 6, under Foster-Lyapunov conditions, we derive explicit error bounds which decay to zero as the quantization gets finer and show better performance than uniform quantization in MDPs with non-compact state spaces under discounted cost criteria. We extend our analysis to the average cost criterion by obtaining explicit error bounds in Theorem 8. (iv) In Theorem 11, we extend our analysis on quantizer design to quantized Q-learning and empirical model learning (which are equivalent in performance). We obtain the error bounds which depend, unlike the planning problem above, on the invariant measure induced by the exploration policy in Theorem 10. For non-compact spaces, in Theorem 11, we show that the approximation error diminishes as quantization becomes finer under Foster–Lyapunov conditions, despite constraints on the dependence of the weighting measures on the exploration policy and quantization bins, and we derive explicit error bounds.
@@ 2.2 Refined Error Bounds for the Discounted Cost Criterion
The above ensures that optimal policies exist. Furthermore, by [23, Lemma 3.19] and [23, Theorem 3.16] (see also [25]), any MDP with a weakly continuous transition probability kernel can be approximated by an MDP with finite action spaces. Accordingly, in the sequel, we assume that $\mathds{U}$ is finite.
@@
where the probability measure $\mathbb{P}_{x_{0}}^{\gamma_{s}}$ over the state and action process is defined by the initial condition $x_{0}$, the policy $\gamma_{s}$, and the kernel $\mathcal{T}$. For a product set $A\times\mathds{U}$ with $A\subseteq\mathds{X}$ measurable we obtain
@@
The following is a supporting result.
@@
Let $X=(X_{1},X_{2},\dots,X_{n})$ be a $\mathds{R}^{n}$-valued random vector with finite expected absolute deviations in each coordinate. Then, the point $y_{i}^{*}=(y_{i,1}^{*},y_{i,2}^{*},\dots,y_{i,n}^{*})$ that minimizes $\mathbb{E}[\|X-y_{i}\|_{1}]$ is obtained by choosing each $y_{i,k}^{*}$ to be a median of the marginal distribution of $X_{k}$ over the bin $B_{i}$, for $k=1,2,\dots,n$.
The case with $n=1$ is proven in [5]. The generalization for the $n$-dimensional case is then immediate: Let $y_{i}\in\mathbb{R}^{n}$ . The expected $\ell_{1}$ distortion writes as:
@@
where $\mu_{k}$ is the marginal distribution of $X_{k}$ over the bin $B_{i}$. We thus can minimize each term separately and the result follows.
@@
Let $\mathds{X}\subseteq\mathbb{R}^{n}$, $b\geq 0$, $\alpha>0$, and define the Lyapunov function $V(x)=\|x\|_{1}^{m}.$ Assume the controlled process $\{X_{t}\}$ satisfies the drift condition
$\mathbb{E}\bigl[V(X_{t+1})\mid X_{t}=x,U_{t}=u\bigr]\leq V(x)-\alpha V(x)+b,\qquad x\in\mathds{X},u\in\mathds{U}$
Under Assumption 2.1 and Assumption 2.2, let $M$ be the total number of hyper-cubic bins in the uniform quantizer. Then, for every initial state $x_{0}\in\mathds{X}$ we have that,
$\bigl|\hat{J}_{\beta}(x_{0})-J^{*}_{\beta}(x_{0})\bigr|\leq\left(\alpha_{c}+\frac{\beta\alpha_{T}\|c\|_{\infty}}{1-\beta}\right)\frac{(2n+1)C^{1/m}}{(M^{\frac{1}{n}(1-\frac{1}{m})})(1-\beta)},$ (8)
where
$C:=\frac{\|x_{0}\|_{1}^{m}(1-\beta)+b\beta}{1-\beta(1-\alpha)}.$
@@ 2.3 Refined Error Bounds for the Average Cost Criterion
The proof is in Appendix D
@@
Assume $\mathds{X}\subseteq\mathbb{R}^{n}$, $f(x)=\|x\|_{1}^{m}$ with $m>1$, and suppose
$\sup_{x,u}\mathbb{E}[V(X_{t+1})\mid X_{t}=x,U_{t}=u]\leq V(x)-f(x)+b,$ (11)
@@
$|\hat{J}_{avg}(x_{0})-J_{avg}^{*}(x_{0})|\leq\left(\alpha_{c}+\frac{\alpha_{T}\|c\|_{\infty}}{\mu(\mathds{X})}\right)\frac{(2n+1)b^{1/m}}{M^{1/n(1-1/m)}}.$ (12)
@@ 3.1 Quantized Q-Learning Algorithm and its Convergence
As noted in [13], quantizing the state space of a continuous MDP converts the problem into a Partially Observable Markov Decision Process (POMDP) which is then non-Markovian. Let $q$ be a quantizer mapping $\mathbb{X}$ to a finite set as described in Section 2.1.
@@
We note that a sufficient condition for the second item above is that the state process $\{X_{t}\}_{t\geq 0}$ is positive Harris recurrent. We have the following convergence result:
Theorem 9 (Theorem 9 [15]).
@@ 3.2 Empirical Model Learning and Equivalence with Quantized Q-Learning
Observe that the above is, see e.g. [14, Theorem 2.1], equal to the almost sure limit of the empirical expression on the right hand side below:
@@
Likewise, by [14, Theorem 2.1], the above is the almost sure empirical limit of of the right hand side below:
@@
An interpretation of the above result then is that one can first obtain the approximate model given with (19-21) by forcing the data into a Markovian model for both the empirical cost estimate (20) and empirical transition kernel estimate (22), and then solve the MDP as if this empirically constructed model is the actual one, instead of running Q-learning whose limit is then optimal precisely for this learned/empirically constructed model. A benefit of such a model-based approach is that one can have better sample complexity bounds compared with Q-learning for certain applications, see e.g. [34] (see also [12, Section 5.1]).
@@
In the planning framework presented in the previous section, we had the flexibility to select the weighting measures $\pi_{y_{i}}$ over the quantization bins arbitrarily. This allowed us to minimize the expected loss by choosing $\pi_{y_{i}}$ as a Dirac measure centered at the median of each bin. In contrast, within the learning context, the weighting measure $\pi_{y_{i}}$ is dictated by the exploration policy and is inherently dependent on the structure of the quantization bins $\{B_{i}\}$.